OpenStudy (anonymous):

Choose the value of the x determinant (Dx) in the following system. 3x – 4y = –7 x + y = 7

OpenStudy (anonymous):

3x-4(7-x)=-7 3x+x-28=-7 4x=28-7=21 x=21/4

OpenStudy (anonymous):

Thank You so much!